6 Suppose that X and Y are random variables such that Var(X)-Var(Y)-2 and Cov(x,y)- 1. Find the value of Var(3.X-Y + 2)
Answer #1

By using the Formula:

Var(aX - bY + c) = Var(aX) + Var(bY) - Cov(aX, bY)

Var(3X - Y + 2)

= Var(3X) + Var(Y) - Cov(3X, Y)

= 32 * Var(X) + Var(Y) - 3 * Cov(X, Y)

= 9*2 + 2 - 3*1

= 18 + 2 - 3

= 17